# Dual Conditioned Motion Diffusion for Pose-Based Video Anomaly Detection

### v1

@@ Introduction
Video Anomaly Detection (VAD) is an essential topic in computer vision and security applications. Different from human action understanding (2018; 2023; 2024), VAD enables prompt identification of abnormal occurrences, such as human actions, accidents, and illnesses. Anomalies are generally characterized as uncommon, unexpected, or unusual phenomena that exhibit significant deviations from normality. Conversely, normality is defined as what is expected and regularly encountered. Detecting video anomalies can be challenging as these events occur infrequently and belong to various categories. Labeling data is costly and time-consuming, making collecting all possible abnormal samples impractical for fully supervised learning approaches. Consequently, anomaly detection problems are typically treated as One Class Classification (OCC) (2020). Only normal data is utilized during model training, and data that deviates significantly from the normal pattern is identified as an anomaly during testing (2021).
There has been a growing trend toward detecting abnormal events or behaviors by analyzing human poses extracted from video frames (Flaborea et al. 2023), as illustrated in Figure 1. Utilizing skeletons to depict human motions in videos is a highly effective approach to protecting privacy and circumventing the limitations of appearance-based attributes. Besides privacy protection, skeletal features are compact, well-structured, and highly descriptive of human motion (Morais et al. 2019).
Existing VAD works are mainly categorized into reconstruction-based and prediction-based methods. Reconstruction-based methods (Li, Chang, and Liu 2021; Chang et al. 2020) first compress the input data into a low-dimensional representation and then recover the original data from this representation. By comparing the reconstruction error between the original data and the recovered data, anomalies such as irregular patterns or structures in the data can be detected. Prediction-based methods (Liu et al. 2017) focus on predicting future frames or events based on history data, and anomalies are indicated when future data deviates from the predicted trends. These methods are effective in modeling and uncovering the temporal connections between successive frames. Nevertheless, they may be prone to noise (Tang et al. 2020).
@@
To address the above issues, we present a unified framework that seamlessly combines the advantages of both reconstruction-based and prediction-based approaches for pose-based VAD. The reconstruction is based on an auto-encoder architecture, whereas the prediction utilizes a diffusion model. We propose a novel Dual Conditioned Motion Diffusion (DCMD) that seamlessly integrates conditioned motion and conditioned embedding, enabling the exploitation of both pose characteristics and latent semantics of observed movements. During the reverse diffusion process, we introduce a motion transformer specifically designed to extract potential correlations from multi-layered spectrum features of human motion. In addition to reconstruction and prediction losses, we devise a United Association Discrepancy (UAD) regularization that leverages Gaussian kernel-based time association and self-attention-based global association. Furthermore, during the inference stage, we employ a mask completion strategy to bolster the utilization of conditioned motion for prediction-based anomaly detection. Experiments conducted on popular human-related anomaly detection datasets, namely HR-ShanghaiTech Campus, HR-Avenue, HR-UBnormal, and UBnormal, demonstrate the superior performance of our proposed method compared to other state-of-the-art approaches.
@@
$\mathcal{T}=\text{Rescale}\left(\left[\frac{1}{\sqrt{2\pi}\sigma_{i}}\mathrm{exp}\left(-\frac{\left|j-i\right|^{2}}{2\sigma_{i}^{2}}\right)\right]_{i,j\in\left\{1,\cdots,(H+F)\right\}}\right),$ (8)
where $\sigma\in{{\mathbb{R}}^{(H+F)\times{h}}}$ denotes the learnable scale parameter for $h$ heads. To transform the association weights into a discrete distribution, we utilize $\mathrm{Rescale}(\cdot)$ by dividing the row sum.
@@ Appendix A Acknowledgments
This work was supported by National Science Foundation of China (62172090, 62302093), Jiangsu Province Natural Science Fund (BK20230833), Start-up Research Fund of Southeast University (RF1028623097), and Big Data Computing Center of Southeast University.

### v2

@@ Introduction
Video Anomaly Detection (VAD) is an essential topic in computer vision and security applications. Different from human action understanding (2018; 2023; 2024; 2025), VAD enables prompt identification of abnormal occurrences, such as human actions, accidents, and illnesses. Anomalies are generally characterized as uncommon, unexpected, or unusual phenomena that exhibit significant deviations from normality. Conversely, normality is defined as what is expected and regularly encountered. Detecting video anomalies can be challenging as these events occur infrequently and belong to various categories. Labeling data is costly and time-consuming, making collecting all possible abnormal samples impractical for fully supervised learning approaches. Consequently, anomaly detection problems are typically treated as One Class Classification (OCC) (2020). Only normal data is utilized during model training, and data that deviates significantly from the normal pattern is identified as an anomaly during testing (2021).
There has been a growing trend toward detecting abnormal events or behaviors by analyzing human poses extracted from video frames (2023), as illustrated in Figure 1. Utilizing skeletons to depict human motions in videos is a highly effective approach to protecting privacy and circumventing the limitations of appearance-based attributes. Besides privacy protection, skeletal features are compact, well-structured, and highly descriptive of human motion (2019).
Existing VAD works are mainly categorized into reconstruction-based and prediction-based methods. Reconstruction-based methods (2021; 2020) first compress the input data into a low-dimensional representation and then recover the original data from this representation. By comparing the reconstruction error between the original data and the recovered data, anomalies such as irregular patterns or structures in the data can be detected. Prediction-based methods (2017) focus on predicting future frames or events based on history data, and anomalies are indicated when future data deviates from the predicted trends. These methods are effective in modeling and uncovering the temporal connections between successive frames. Nevertheless, they may be prone to noise (2020).
@@
To address the above issues, we present a unified framework that seamlessly combines the advantages of both reconstruction-based and prediction-based approaches for pose-based VAD. The reconstruction is based on an auto-encoder architecture, whereas the prediction utilizes a diffusion model. We propose a novel Dual Conditioned Motion Diffusion (DCMD) that seamlessly integrates conditioned motion and conditioned embedding, enabling the exploitation of both pose characteristics and latent semantics of observed movements. During the reverse diffusion process, we introduce a motion transformer specifically designed to extract potential correlations from multi-layered spectrum features of human motion. In addition to reconstruction and prediction losses, we devise a United Association Discrepancy (UAD) regularization that leverages Gaussian kernel-based time association and self-attention-based global association. Furthermore, during the inference stage, we employ a mask completion strategy to bolster the utilization of conditioned motion for prediction-based anomaly detection. Experiments conducted on popular human-related anomaly detection datasets demonstrate the superior performance of our proposed method compared to state-of-the-art approaches.
@@
$\mathcal{T}=\text{Rescale}\left(\left[\frac{1}{\sqrt{2\pi}\sigma_{i}}\mathrm{exp}\left(-\frac{\left|j-i\right|^{2}}{2\sigma_{i}^{2}}\right)\right]\right),$ (8)
where $\sigma\in{{\mathbb{R}}^{(H+F)\times{h}}}$ denotes the learnable scale parameter for $h$ heads and $i,j\in\left\{1,\cdots,(H+F)\right\}$. To transform the association weights into a discrete distribution, we utilize $\mathrm{Rescale}(\cdot)$ by dividing the row sum.
@@ Appendix A Acknowledgments
This work was supported by National Science Foundation of China (62172090, 62302093, 52441503), Jiangsu Province Natural Science Fund (BK20230833), Start-up Research Fund of Southeast University (RF1028623097), and Big Data Computing Center of Southeast University.